if the extremities of a diagonal of a square are (1,2,3) and (2,-3,5) then the length of it’s side is?

Grade:12th Pass

1 Answers

Shahid
7 years ago
Dear Charitha ,
I want you to note that it’s a Cube , not a square which is formed by those points.(Because it’s in 3D Geometry)
Now, 
Given:- Let A (1,2,3) and E (2,-3,5) be the Extrimities of the given “Cube”.(say,ABCDEFGH)
To Find:- The Length of the side of the cube.
Solution: 
First, let us Find the length AE.
AE = \sqrt{(2 - 1)^{2} + (2 - (-3))^{2} + (3 - 5)^{2}))}\sqrt{}
AE = \sqrt{30}
 
Here , AE is the Body Diagonal.
And Let AB be a Side of the cube.
Let length of AB = a
 
Now By Solving , We get that AE = \sqrt{3} x a
\therefore AE = 3 \sqrt{10}